You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked youLydex Company Comparative Income Statement and Reconciliation This Year Last Year Sales (all on account) $ 15,940,000 $ 14,38Required: 1. You decide first to assess the companys performance in terms of debt management and profitability. Compute the
Answer #1
a)
Times Interest Earned Ratio = EBIT/ Interest expenses
This Year Last Year
= $1972000/372000 = $1559000/312000
=5.3 times =5 times
b) Debt To Equity Ratio = Total Debt / Total Equity
This Year Last Year
= $7810000/9880000 = $6260000/9080000
=0.79 times =0.69 times
c)
Gross Profit Margin = Gross Profit / Sales
This Year Last Year
= $3188000/15940000 = $3595000/14380000
=20 % =25 %
d) Average Assets = (beginning Assets + ending Assets)/2
This Year Last Year
= ( $15340000+17690000)/2 = ( $13150000+15340000)/2
= $ 16515000 = $ 14245000
Return On Assets = Net Income / Average Assets
This Year Last Year
= $1120000/16515000 = $1152900/14245000
=6.8% =8.1 %
e) Average Stockholders Equity = (beginning Stockholders Equity + ending Stockholders Equity)/2
This Year Last Year
= ( $9080000+9880000)/2 = ( $8503350+9080000)/2
= $ 9480000 = $ 8791675
Return On Stockholders Equity = Net Income / Average Stockholders Equity
This Year Last Year
= $1120000/9480000 = $1152900/8791675
=11.8 % =13.1 %
f) Positive Positive
Let me know the wrong answer. I will correct immediately. Please vote positively.